Multigrain Health Foods Inc. is authorized to issue 5,000,000 common shares. In its initial public offering during 2010, Multigrain issued 500,000 common shares for $7.00 per share. Over the next year, Multigrain’s share price increased and the company issued 400,000 more shares at an average price of $8.50.

During the next seven years, from 2010 to 2016, Multigrain earned net income of $920,000. They declared and paid cash dividends of $140,000. A 10% stock dividend was distributed to the shareholders in 2016 on the shares outstanding. The market price was $8.00 per share when the stock dividend was distributed. At December 31, 2016, the company has total assets of $14,500,000 and total liabilities of $6,820,000.

Show the computation of Multigrain’s total shareholders’ equity at December 31, 2016. Present a detailed computation of each element of shareholder’s equity.
Answer #1

Answer:

Step 1:Computation of Number of shares, Stock dividend,Value of stock dividend.

a.Number of shares =500,000 + 400,000 = 900,000 Shares.

b.Stock dividend = 10% * 900,000 = 90,000 Shares

c.Value of stock dividend = 90000 * 8 = $720,000

Step 2:Computation of Multigrain’s Total shareholders’ equity at December 31, 2016.

Particulars Amount($)
Common stock Capital
500,000 shares @ $7 each 3,500,000.00
400,000 shares @ $8.5 each 3,400,000.00
Add:Stock Dividend 90000 shares@ $8 each 720,000.00
A = 7,620,000.00
Multigrain Retained Earnings 2010 to 2016 Amount($)
Net income 920,000.00
Less: Cash Dividend 140,000.00
= 780,000.00
Less: Stock dividend(90000 shares@ $8 each) 720,000.00
B =   60,000.00
Multigrain’s Total shareholders’ equity at December 31, 2016=(A+B)= 7,680,000.00